OpenStudy (anonymous):

Find the slope of the line that contains the following pair of points. (4,3) , (6,7)

OpenStudy (judygreeneyes):

slope = (y2 - y1) / (x2 - x1) = (7 - 3)/(6 - 4) = 4/2 = 2

OpenStudy (anonymous):

it positive 2???

OpenStudy (judygreeneyes):

yes, positive 2